OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x))

OpenStudy (anonymous):

f(g(x)) = f(10x-6) = -4(10x-6)+7 = (-40x+24)+7 = -40x+31